OpenStudy (anonymous):

I need help from anyone who understands exponential growth functions! Pretty please :))

OpenStudy (anonymous):

whats your question?

OpenStudy (anonymous):

tag me when your done posting it

OpenStudy (anonymous):

I need to solve for "t" I don't want to put the actual question, but here's an example. \[200=5000(1+.015)^(12t)\]

OpenStudy (anonymous):

t = 0.00328407

OpenStudy (anonymous):

can you give me a medal for helping please @brianakf

OpenStudy (anonymous):

I want you to show me how you got that

OpenStudy (anonymous):

I don't know how to do it

OpenStudy (anonymous):

alright but is this a test ?

OpenStudy (anonymous):

\\[200=5000(1+.015)^{12t}\] seems odd

OpenStudy (anonymous):

no, homework

OpenStudy (anonymous):

what was the actual question?

OpenStudy (anonymous):

first off \(1+.015=1.015\) so this is \[200=5000(1.015)^{12t}\] but it still seems strange

OpenStudy (anonymous):

200 is A (amount), 5000 P (principle) 1.5% r (rate). T (time)

OpenStudy (anonymous):

Its the same, but I changed the numbers. The teacher has something that checks online to make sure we don't use these websites

OpenStudy (anonymous):

how can you invest $5000 and earn interest and end up with only $200

OpenStudy (anonymous):

No, its a decay. You owe 5000, you want to know how long until you only owe 200

OpenStudy (anonymous):

then this is not the formula to use you need to know the payoff formula but if you are supposed to solve \[200=5000(1.015)^{12t}\] you can do it in two steps

OpenStudy (anonymous):

I'm supposed to use that formula

OpenStudy (anonymous):

sorry i lost connection but i guess ^ can help you

OpenStudy (anonymous):

@shinebrightlikeadimon if he can't ill tag you so you can!

OpenStudy (anonymous):

divide by \(5000\) and get \[(1.015)^{12t}=.04\] then use the change of base formula and get \[12t=\frac{\ln(.04)}{\ln(1.015)}\]

OpenStudy (anonymous):

making \[t=\frac{\ln(.04)}{12\ln(1.015)}\]

OpenStudy (anonymous):

but this just solves the equation you wrote, not at all the question that was asked

OpenStudy (anonymous):

How would I get an actual number for t though, is that even possible?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I just solved it thanks!

OpenStudy (anonymous):

wait was my answer correct

OpenStudy (anonymous):

no it was around -18

OpenStudy (anonymous):

The answer is supposed to be negative

OpenStudy (anonymous):

sorry i did my best

OpenStudy (anonymous):

it's ok, how do you give medals?

OpenStudy (anonymous):

you click on best response and the person gets the medal